OpenStudy (lukecrayonz):

Use a graphing utility to graph the function, then use the table feature to find the limit.

OpenStudy (lukecrayonz):

\[\lim_{x \rightarrow 5} \lim(\ln(x+3))\]

OpenStudy (lukecrayonz):

@satellite73 any help for this? Do I just go to table and put 5, 4.999, 4.99, 4.9 etc?

OpenStudy (anonymous):

yes if that is what it asks

OpenStudy (lukecrayonz):

Well I'm not sure if thats what it's asking, just making sure.

OpenStudy (anonymous):

i assume it means \[\lim_{x\to 5}\ln(x+3)\] which is clearly \[\ln(8)\] because log is a continuous function

OpenStudy (lukecrayonz):

Why ln(8)?

OpenStudy (anonymous):

because \(5+3=8\)

OpenStudy (lukecrayonz):

oh...

OpenStudy (anonymous):

\[\lim_{x\to 5}\ln(x+3)=\ln(\lim_{x\to 5}x+3)=\ln(5+3)=\ln(8)\]

OpenStudy (lukecrayonz):

Well i just found out its asking for me to create a table like these questions: http://screensnapr.com/v/7oD9V1.png

OpenStudy (anonymous):

ok you can do it it is silly, but you can do it replace x by values close to 5

OpenStudy (lukecrayonz):

Well yes, but thats what the table function is just doing for you. Ti-84, 2nd Graph, and it'll show all the values, and you can input values, but im sure you know that.